A 32.1 L metal cylinder at 29°C is filled with an ideal gas at 15.2 ATM. What is the pressure in the cylinder if it is heated to 105°C
1
Expert's answer
2020-06-09T07:22:11-0400

As V = const

P1/T1 = P2/T2

15.2 / (29+273.15) = P2 / (105+273.15)

P2 = 19.02 ATM